HUBBLE SPACE TELESCOPE DAILY REPORT #2149 PERIOD COVERED: 0000Z (UTC) 06/19/98 - 0000Z (UTC) 06/22/98 Daily Status Report as of 173/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Seven Sets of NIC/1/2/3 7817 (A Complete NICMOS Map of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations to map the complete WF/PC-2 region of the central Hubble Deep Field. These data will provide unique information about the faintest objects, but will include only a small number of objects with spectroscopic redshifts or z ~ 3 Lyman break galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Seven Sets of WF/PC-2 8013 (WF/PC-2 Parallel Images to the HDF Flanking Fields {Companion Proposal to 7817}) The WF/PC-2 was used to make observations parallel to the Hubble Deep Field observations (i.e., flanking fields). The complete WF/PC-2 region of the central Hubble Deep Field {HDF} was being mapped with the NICMOS Camera 3 at F110W and F160W {proposal 7817}. While these observations were executing, the WF/PC-2 was working in parallel mode to image areas near this field. The observations were completed as scheduled, and no anomalies were reported. 1.4 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.5 Completed Three Sets of NIC/1/2/3 7822 (Grism Spectra of Centaurs and TNOs from 1-2 Microns) The Near Infrared Camera (Cameras One, Two and Three) was used to make grism spectra of objects Chiron, 1993HA2 and 1996RR20 from 1-2 microns. The surfaces of Centaurs and Trans-Neptunian Objects (TNOs) contain potentially unique records of early evolutionary processing {accretional and dynamical} in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Five Sets of NIC/3 7830 (IR Spectra for Known-Mass M Dwarfs) The Near Infrared Camera (Camera Three) was used to make infrared spectrographic observations of the M-Class dwarf stars KELU1, GJ22ABC, GJ699, GJ752B, GJ1245ABC,GJ860AB, GJ1005AB and GJ406. Over the last decade there has been considerable debate on two major issues in the low-mass star literature: {1} the effective temperature scale for M dwarfs and its relation to mass and luminosity and {2} the conversion from an absolute magnitude in some passband to a bolometric magnitude. During the first iteration of this proposal, the default to fine lock detailed in 2.1 occurred. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.7 Completed Three Sets of NIC/1/2/3 7322 (The IMF Below 1 Msun in Young Clusters: 1.9 Micron Water Band Imaging of IC348) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of clusters GL-699, BD+01D4774, VB-10. This study will be the first to obtain spectral types for a young cluster population complete to K=16, and thereby quantify, to unprecedentedly high precision, the IMF below 1M{sun} in young clusters. The observations were executed as scheduled, and no anomalies were reported. 1.8 Completed Seven Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.9 Completed Three Sets of NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of the galaxies NGC-151, NGC-2998 and NGC-6699 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.10 Completed Two Sets of NIC/1/2/3 7456 (Water in Protoplanetary Disks) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of the protoplanetary disk around stars 1548C27 and AS353A to search for water in their disks. This survey will be the first to take a detailed look at the gas phase properties of pre-main-sequence disks at AU distances. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed NIC/3 7841 (HST/NICMOS Observations of the Nearest Embedded Cluster: the L1688 {Rho Oph} Protostellar Region) The Near Infrared Camera (Camera Three) was used to make observations protostellar region of the nearest young cluster L1688 (Positions 5,6 and 7) ( a.k.a., Rho Oph). This cluster is unique in its proximity to Earth, only 125 pc away; and in its high proportion of optically invisible protostars. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed NIC/3 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera Three) was used to make narrow band observations of the BL-Lac (BAL) QSO Q1029-0125. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.14 Completed NIC/3 7816 (NIC/3 Photometric Zero Point) The Near Infrared Camera (Camera Three) was used to make observations of standard star P330E-CAL in order to study the zero point flux for a set of filters. The NIC/3 intra-pixel sensitivity will also be measured. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 31 Successful : 31 Scheduled Reacquisitions : 18 Successful : 18 HSTAR 65647 documents that the guide star acquisition at 170/122907Z defaulted to fine lock backup on FGS-3 when the scan step limit on FGS-1 was exceeded. The proposal detailed in 1.6 may have been affected. 2.2 FHST Updates: Scheduled : 54 Successful : 54 2.3 Operations Notes: Per operations request, real-time FHST maps were performed four times. Using ROP SR-1, the SSR EDAC error counter was cleared five times. Per Operations Note 446, the DF-224 was successfully dumped to the CCS A and B strings at 171/0452Z, 171/1526Z, 172/0336Z, 172/2002Z, and 173/0122Z. The ground master image was updated at 172/0411Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH